find two numbers such that the sum of thrice the first and the second is 142 and four times the first exceeds the second by 138
Answer by dkppathak(439)

let number is x and y
3x+y=142
4x-y =138
by solving elimination method
7x=280
x=280/7
x=40
3x+y=142
120+y=142
y=142-120
y =22
ANSWER
first number is 40
and second number is 22